# Causal Inference in Longitudinal Data under Unknown Interference

### v4

@@ 2.2 Estimands
In the example introduced previously, we can define a series of AMRs for each value of $\mathbf{a}^{(1:T)}$ and $d\in\mathcal{D}=\{0,1,\dots,7\}$, where $d$ is measured by the shortest length path in the social network and the Euclidean distance in the spatial raster. $\Omega_{j}(d)$ is defined as $j$’s $d$th-degree neighbors in the former case and a doughnut surrounding $j$ with $\kappa=1$ in the latter, as depicted in the top plots of Figure 2. The top row of Table 2 presents AMRs defined under each of the four possible treatment histories, which can be summarized by the following MSM at any $d$:
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\beta_{0}(d)+\beta_{1}(d)a^{(1)}+\beta_{2}(d)a^{(2)}+\beta_{3}(d)a^{(1)}a^{(2)},\vspace{-1em}$
@@
We specify an effect function that emanates from each unit and declines in $d$ exponentially. Effects received by each unit are additive across other units in the sample and amplified by an idiosyncratic constant that represents treatment effect heterogeneity (depicted by varying colors in the plot). The effects in each period carry over into the next, declining by $50\%$ in magnitude. In period $t$, the probability for unit $i$ to receive the treatment depends on its treatment status and the value of a time-varying confounder in the previous period. We repeat the simulation $1,000$ times and approximate each $Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ by the average of $Y_{i}\left(\mathbf{a}^{(1:T)};\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ over the simulations. The AMRs are then constructed in line with their definition in Equation (4). We plot several linear combinations of the MSM parameters against proximity levels in the bottom-right plot of Figure 2. As shown, each curve captures how the effects generated by a specific treatment history, relative to the benchmark $(0,0)$, vary with proximity.
@@ 3.1 Identification Assumptions
$\Big{\{}Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right),\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\Big{\}}\perp A_{j}^{(t)}\mid\left(\mathbf{A}_{j}^{(1:(t-1))},\mathbf{L}_{j}^{(1:t)}\right).\vspace{-1em}$
@@
Assumption 1 states that for any unit $j$, its treatment status in period $t$, $A_{j}^{(t)}$, is independent to the distribution of any unit’s potential outcomes or the potential values of time-varying confounders in period $t+1$—rather than just its own—conditional on its history of treatment assignment and time-varying confounders. The assumption reduces to the familiar version in the absence of interference, where $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=Y_{i}\left(\mathbf{a}^{(1:T)}\right)$, $\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)}\right)$, and units are independent from each other. Assumption 1 is violated if the treatment assignment of unit $j$ depends on the history of another unit, thereby excluding any form of diffusion or dependence in treatment across units (VanderWeele and An, 2013; Sävje, Aronow and Hudgens, 2021). This requirement is naturally satisfied in experiments and observational studies where a unit’s treatment status is determined independently, without reference to others, and commonly adopted in the literature (e.g., Liu, Hudgens and Becker-Dreps, 2016). In Section 5, we explore ways to extend the method to settings where treatment diffusion exists or treatment assignment is correlated across units within the same period.
Assumption 1 tends to be weaker than the assumptions required for exposure mapping methods, which demand that the level of exposure for any unit $i$—typically a function of treatment histories within a subset of $i$’s neighbors—is independent of $i$’s potential outcomes conditional on its history (Forastiere, Airoldi and Mealli, 2021). This requirement fails in cases where $Y_{i}$ is influenced by $\mathbf{L}_{j}^{(T-1)}$, a phenomenon known as “contextual factors” when $\mathbf{L}_{j}^{(T-1)}$ contains time-invariant variables (VanderWeele and An, 2013; Egami and Tchetgen Tchetgen, 2024) or “contagion” when it incorporates past values of the outcome (Ogburn and VanderWeele, 2017). In contrast, Assumption 1 remains valid even in these scenarios.
@@
$1-\eta<P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)<\eta.\vspace{-0.5em}$
@@
Next, we impose the restriction that any time-varying confounder is influenced only by treatment histories up to the present and not by future treatment statuses. This assumption, commonly referred to as no anticipation (e.g., Hernán and Robins, 2010), is formally stated as follows:
@@
$\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)}\right)=\mathbf{L}_{i}^{(t+1)}\left(\tilde{\mathbf{a}}^{(1:T)}\right)\text{ if }\mathbf{a}^{(1:t)}=\tilde{\mathbf{a}}^{(1:t)}.\vspace{-1em}$
Although Assumption 3 is not essential for causal identification, it simplifies estimation and facilitates placebo tests (see Section 5 for details). In Section 5, we discuss possible relaxations of this assumption. We present structural equations that describe the relationships among variables under Assumptions 13 and discuss their connection to the framework in Ogburn et al. (2020) in the Supplementary Material.
@@ 3.2 Estimator
The key insight of the paper is that assumptions introduced in the previous subsection are sufficient for identifying the AMR at any $d$ and the parameters in the associated MSM. Define the transformed outcome of unit $j$ at $d$ as
$\mu_{j}(d)=\mu\left(\big{\{}Y_{i}\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)=\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i}}{|\Omega_{j}(d)|},\vspace{-1em}$ (6)
@@
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right].\vspace{-1em}$ (7)
Since each parameter in the MSM is a linear combination of the AMRs, they can be similarly identified from the data. The proofs are provided in the Supplementary Material.
@@
$\hat{\boldsymbol{\beta}}(d)={\rm arg}\,\min_{\boldsymbol{\beta}(d)}\sum_{j=1}^{N}w_{j}\left(\mu_{j}(d)-\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}(d)\right)^{2},\vspace{-1em}$ (8)
where $w_{j}=\frac{\prod_{s=1}^{T}P\left(A_{j}^{(s)}\mid\mathbf{A}_{j}^{(1:(s-1))}\right)}{e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}$ is the stabilized weight (Cole and Hernán, 2008). If $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ is unknown, as in an observational study, then we replace it with an estimate in $w_{j}$. In the following section, we show that this estimator possesses desirable large-sample properties, such as consistency and asymptotic normality, under mild restrictions on the dependence across units caused by interference.
Beyond its validity under unknown interference, the proposed method offers two other advantages over approaches based on exposure mapping. First, it only requires researchers to estimate the probability of receiving treatment, rather than the probability of attaining a specific level of exposure, which often involves computing a convolution. [^7] For instance, in a common exposure mapping where exposure is defined as the average of treatment statuses among a unit’s first-degree neighbors, computing this probability is equivalent to deriving the distribution of the sum of multiple random variables. Moreover, the exposure probabilities can be close to 0 or 1 even when Assumption 2 is satisfied. Second, because identification relies on Assumption 1, $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ does not need to account for variables influencing the treatment statuses of a unit’s neighbors, such as contextual factors or sources of “spatial confounding” (Papadogeorgou, Choirat and Zigler, 2019; Reich et al., 2021). Steps 1-4 in Algorithm 1 summarize the estimation procedure for the proposed method. We provide justifications for Steps 4 and 5 in the following section.
@@ 4 Large-Sample Theory
To derive the large-sample properties of our estimator, we consider a sequence $\{\mathcal{G}_{N}\}_{N>0}$ in which $\mathcal{N}_{N}$ units are embedded, with $N\to\infty$. Define the vector of transformed outcomes at proximity level $d$: $\boldsymbol{\mu}(d)=(\mu_{1}(d),\mu_{2}(d),\dots,\mu_{N}(d))^{\prime}$, the vector of outcomes: $\mathbf{Y}=(Y_{1},Y_{2},\dots,Y_{N})^{\prime}$, and the $N\times N$ matrix: $\mathbf{D}(d)=\Big{\{}\frac{\mathbf{1}\{i\in\Omega_{j}(d)\}}{|\Omega_{j}(d)|}\Big{\}}_{N\times N}$.
From Equation (6), we can see that $\boldsymbol{\mu}(d)=\mathbf{D}(d)\mathbf{Y}$. Next, define the matrix of regressors for the MSM as: $\mathbf{M}=\left(\mathbf{m}\left(\mathbf{A}_{1}^{(1:T)}\right),\mathbf{m}\left(\mathbf{A}_{2}^{(1:T)}\right),\dots,\mathbf{m}\left(\mathbf{A}_{N}^{(1:T)}\right)\right)^{\prime}$, and the diagonal weight matrix: $\mathbf{W}=diag\left(w_{1},w_{2},\dots,w_{N}\right)$. We know that estimates from the WLS algorithm can be expressed as
@@
$=$ $\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\mathbf{Z}^{\prime}(d)\mathbf{Y}\right)=\left(\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\right),\vspace{-1em}$
@@
Since $\mathbf{A}_{i}^{(1:T)}$ is conditionally independent across units, we can show the convergence of the first term, $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$, under standard regularity conditions. For the second term, we know that $\mathbf{Z}_{i}(d)=\sum_{j=1}^{N}\frac{\mathbf{1}\{j\in\Omega_{i}(d)\}\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right)w_{j}}{|\Omega_{j}(d)|}$, which is a weighted average of the MSM regressors across $i$’s neighbors in $\Omega_{i}(d)$. $\mathbf{Z}_{i}(d)$ and $\mathbf{Z}_{j}(d)$ are independent if $\Omega_{i}(d)$ and $\Omega_{j}(d)$ do not overlap. [^8] Suppose $\Omega_{j}(d)\coloneqq\{i\in\mathcal{N}:d_{ij}=d\}$, then this is true when $d_{ij}>2d$. Therefore, the large-sample performance of $\hat{\boldsymbol{\beta}}(d)$ is primarily driven by the dependence in $Y_{i}$, which may be caused by either interference or correlated contextual factors.

### v5

@@ 2.2 Estimands
In the example introduced previously, we can define a series of AMRs for each value of $\mathbf{a}^{(1:T)}$ and $d\in\mathcal{D}=\{0,1,\dots,7\}$, where $d$ is measured by the length of the shortest path connecting two units in the social network and the Euclidean distance between two units in the spatial raster. $\Omega_{j}(d)$ is defined as $j$’s $d$th-degree neighbors in the former case and a doughnut surrounding $j$ with $\kappa=1$ in the latter, as depicted in the top plots of Figure 2. The top row of Table 2 presents AMRs defined under each of the four possible treatment histories, which can be summarized by the following MSM at any $d$:
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\beta_{0}(d)+\beta_{1}(d)a^{(1)}+\beta_{2}(d)a^{(2)}+\beta_{3}(d)a^{(1)}a^{(2)},\vskip-10.00002pt$
@@
We specify an effect function that emanates from each unit and declines in $d$ exponentially. Effects received by each unit are additive across other units in the sample and amplified by an idiosyncratic constant that represents treatment effect heterogeneity. The effects in each period carry over into the next, declining by $50\%$ in magnitude. In period $t$, the probability for unit $i$ to receive the treatment depends on its treatment status and the value of a time-varying confounder in the previous period. We repeat the simulation $1,000$ times and approximate each $\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ by the average of $Y_{i}\left(\mathbf{a}^{(1:T)};\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ over the simulations. The AMRs are then constructed in line with their definition in Equation (4). We plot several linear combinations of the MSM parameters against proximity levels in the bottom-right plot of Figure 2. As shown, each curve captures how the effects generated by a specific treatment history, relative to the benchmark $(0,0)$, vary with proximity. [^8] The two histories, $(1,0)$ and $(0,1)$, produce spillover effects through different mechanisms. For $(0,1)$, the effect arises from the spillover of contemporaneous treatments. For $(1,0)$, it reflects both the spillover of past treatments and their influence on the distribution of contemporaneous treatments.
@@ 3.1 Identification Assumptions
$\Big{\{}Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right),\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\Big{\}}\perp A_{j}^{(t)}\mid\left(\mathbf{A}_{j}^{(1:(t-1))},\mathbf{L}_{j}^{(1:t)}\right).\vskip-10.00002pt$
@@
Assumption 1 states that for any unit $j$, its treatment status in period $t$, $A_{j}^{(t)}$, is independent to the distribution of any unit’s potential outcomes or the potential values of time-varying confounders in period t+1t+1—rather than just its own—conditional on its history of treatment assignment and time-varying confounders. [^9] Our identification result only requires Assumption 1 to hold for the potential outcomes. However, the broader formulation allows for the possibility of evaluating the assumption’s validity through placebo tests, as discussed in Section 5. The assumption reduces to the familiar version in the absence of interference, where $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=Y_{i}\left(\mathbf{a}^{(1:T)}\right)$, $\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)}\right)$, and units are independent from each other. Assumption 1 is violated if the treatment assignment of unit $j$ depends on the history of another unit, thereby excluding any form of diffusion or dependence in treatment across units (VanderWeele and An, 2013; Sävje, Aronow and Hudgens, 2021). This requirement is naturally satisfied in experiments and observational studies where a unit’s treatment status is determined independently and commonly adopted in the literature (e.g., Liu, Hudgens and Becker-Dreps, 2016). In such cases, Assumption 1 holds even in the presence of “contextual factors,” where the confounders are correlated across units (VanderWeele and An, 2013; Egami and Tchetgen Tchetgen, 2024), [^10] A common example is sources of “spatial confounding” discussed in Papadogeorgou, Choirat and Zigler (2019) and Reich et al. (2021). or when contagion exists and the confounders affect another unit’s outcome (Ogburn and VanderWeele, 2017). In Section 5, we explore ways to extend the method to settings where treatment diffusion exists or treatment assignment is correlated across units.
@@
$1-\eta<P\left(A_{j}^{(t)}=1\mid\mathbf{V}_{j}^{(t)}\right)<\eta.\vskip-10.00002pt$
@@
Next, we impose the restriction that any time-varying confounder is influenced only by treatment histories up to the present and not by future treatment statuses. This assumption, commonly referred to as no anticipation, is formally stated as follows:
@@
$\mathbf{L}_{i}^{(t+1)}\left(\mathbf{a}^{(1:T)}\right)=\mathbf{L}_{i}^{(t+1)}\left(\tilde{\mathbf{a}}^{(1:T)}\right)\text{ if }\mathbf{a}^{(1:t)}=\tilde{\mathbf{a}}^{(1:t)}.$
Although Assumption 3 is not essential for causal identification, it simplifies estimation and facilitates placebo tests (see Section 5 for details). In Section 5, we discuss possible relaxations of this assumption. We present structural equations that describe the relationships among variables under Assumptions 13 and discuss their connection to the framework in Ogburn et al. (2020) in Section A.1 of the Supplementary Material.
@@ 3.2 Estimator
The key insight of the paper is that assumptions introduced in the previous subsection are sufficient for identifying the AMR at any proximity level $d$ and the parameters in the associated MSM. Define the transformed outcome of unit $j$ at $d$ as
$\mu_{j}(d)=\mu\left(\big{\{}Y_{i}\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)=\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i}}{|\Omega_{j}(d)|},\vskip-10.00002pt$ (6)
@@
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\frac{1}{N}\sum_{j=1}^{N}{\rm E}\,\left[\frac{\mathbf{1}\{\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)}\}\mu_{j}(d)}{e\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}\right].\vskip-10.00002pt$ (7)
Since each parameter in the MSM is a linear combination of the AMRs, they can be similarly identified from the data. The proofs are in Section B.1 of the Supplementary Material.
@@
$\hat{\boldsymbol{\beta}}(d)={\rm arg}\,\min_{\boldsymbol{\beta}(d)}\sum_{j=1}^{N}w_{j}\left(\mu_{j}(d)-\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}(d)\right)^{2},\vskip-10.00002pt$ (8)
where $w_{j}=\frac{\prod_{s=1}^{T}P\left(A_{j}^{(s)}\mid\mathbf{A}_{j}^{(1:(s-1))}\right)}{e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)}$ is the stabilized weight (Cole and Hernán, 2008). If $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ is unknown, as in observational studies, then we replace it with an estimate in $w_{j}$. In the following section, we show that this estimator possesses desirable large-sample properties, such as consistency and asymptotic normality, under mild restrictions on the dependence across units caused by interference. When $d=0$, $\mu_{j}(d)=Y_{j}$, and Equation (8) becomes the classic estimator in the absence of interference. Therefore, results from the classic estimator can be seen as estimated effects for the AMR defined at 0.
Beyond its validity under unknown interference, the proposed method offers two other advantages over approaches based on exposure mappings. First, it only requires researchers to estimate the probability of receiving treatment, rather than the probability of attaining a specific level of exposure, which often involves computing a convolution. [^11] For instance, in a common exposure mapping where exposure is defined as the average of treatment statuses among a unit’s first-degree neighbors, computing this probability is equivalent to estimating the distribution of the sum of multiple random variables. Moreover, the exposure probabilities can be close to 0 or 1 even when Assumption 2 is satisfied. Second, because identification relies on Assumption 1, $e\left(\mathbf{A}_{j}^{(1:T)};\mathbf{V}_{j}^{(1:T)}\right)$ does not need to account for variables influencing the treatment statuses of a unit’s neighbors (Forastiere, Airoldi and Mealli, 2021). Steps 1-4 in Algorithm 1 summarize the estimation procedure for the proposed method. We provide justifications for Steps 4 and 5 in the following section.
@@ 4 Large-Sample Theory
To derive the large-sample properties of our estimator, we consider a sequence $\{\mathcal{G}_{N}\}_{N>0}$ in which $\mathcal{N}_{N}$ units are embedded, with $N\to\infty$ and $T$ fixed. Define the vector of transformed outcomes at proximity level $d$: $\boldsymbol{\mu}(d)=(\mu_{1}(d),\mu_{2}(d),\dots,\mu_{N}(d))^{\prime}$, the vector of outcomes: $\mathbf{Y}=(Y_{1},Y_{2},\dots,Y_{N})^{\prime}$, and the $N\times N$ matrix: $\mathbf{D}(d)=\Big{\{}\frac{\mathbf{1}\{i\in\Omega_{j}(d)\}}{|\Omega_{j}(d)|}\Big{\}}_{N\times N}$.
From Equation (6), we can see that $\boldsymbol{\mu}(d)=\mathbf{D}(d)\mathbf{Y}$. Next, define the matrix of regressors for the MSM as: $\mathbf{M}=\left(\mathbf{m}\left(\mathbf{A}_{1}^{(1:T)}\right),\mathbf{m}\left(\mathbf{A}_{2}^{(1:T)}\right),\dots,\mathbf{m}\left(\mathbf{A}_{N}^{(1:T)}\right)\right)^{\prime}$, and the diagonal weight matrix: $\mathbf{W}=diag\left(w_{1},w_{2},\dots,w_{N}\right)$. Estimates from the WLS algorithm can be expressed as
@@
$=$ $\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\mathbf{Z}^{\prime}(d)\mathbf{Y}\right)=\left(\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)\right)^{-1}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{Z}_{i}(d)Y_{i}\right),\vskip-10.00002pt$
@@
Since $\mathbf{A}_{i}^{(1:T)}$ is conditionally independent across units, we can show the convergence of the first term, $\frac{1}{N}\sum_{i=1}^{N}w_{i}\mathbf{m}\left(\mathbf{A}_{i}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{i}^{(1:T)}\right)$, under standard regularity conditions. For the second term, $\mathbf{Z}_{i}(d)=\sum_{j=1}^{N}\frac{\mathbf{1}\{j\in\Omega_{i}(d)\}\mathbf{m}\left(\mathbf{A}_{j}^{(1:T)}\right)w_{j}}{|\Omega_{j}(d)|}$, which is a weighted average of the MSM regressors across $i$’s neighbors in $\Omega_{i}(d)$. $\mathbf{Z}_{i}(d)$ and $\mathbf{Z}_{j}(d)$ are uncorrelated if $\Omega_{i}(d)$ and $\Omega_{j}(d)$ do not overlap. [^12] Suppose $\Omega_{j}(d)\coloneqq\{i\in\mathcal{N}:d_{ij}=d\}$, then this is true when $d_{ij}>2d$. Therefore, the large-sample performance of $\hat{\boldsymbol{\beta}}(d)$ is primarily driven by the dependence in $Y_{i}$, which may be caused by either interference or correlated contextual factors.